OpenStudy (anonymous):

Solve the quadratic equation by the square root property. (2x + 5)^2 = 49

OpenStudy (solomonzelman):

\[a^2=b^2~~~~~~~~~~a=b\]

OpenStudy (solomonzelman):

use this here.

OpenStudy (anonymous):

i did but cant get the right answer i missed class so i have no clue what to do

OpenStudy (anonymous):

I used the quadratic formula and got x=1,−6

OpenStudy (anonymous):

yeah thats what i chose too

OpenStudy (anonymous):

well then you good ha

OpenStudy (solomonzelman):

\[(2x + 5)^2 = 49\]\[(2x + 5)^2 = 7^2\]\[(2x+5)=7\]\[2x=2\]\[x=1\]

OpenStudy (solomonzelman):

\[(2~(1)~+5~)^2=~49\]\[(2+5)^2=49\]\[7^2=49\]\[49=49\]\[CHECKED!\]